# LoRA is All You Need for Safety Alignment of Reasoning LLMs

### v1

@@ Abstract
Reasoning LLMs have demonstrated remarkable breakthroughs in solving complex problems that were previously out of reach. To ensure LLMs do not assist with harmful requests, safety alignment fine-tuning is necessary in the post-training phase. However, safety alignment fine-tuning has recently been shown to significantly degrade reasoning abilities, a phenomenon known as the “Safety Tax”. In this work, we show that using LoRA for SFT on refusal datasets effectively aligns the model for safety without harming its reasoning capabilities. This is because restricting the safety weight updates to a low-rank space minimizes the interference with the reasoning weights. Our extensive experiments across four benchmarks covering math, science, and coding show that this approach produces highly safe LLMs—with safety levels comparable to full-model fine-tuning—without compromising their reasoning abilities. Additionally, we observe that LoRA induces weight updates with smaller overlap with the initial weights compared to full-model fine-tuning. We also explore methods that further reduce such overlap—via regularization or during weight merging—and observe some improvement on certain tasks. We hope this result motivates designing approaches that yield more consistent improvements in the reasoning–safety trade-off.
@@ 1 Introduction
Our extensive experiments reveal the surprising effectiveness of a simple recipe for safety alignment of reasoning models: Applying LoRA during SFT using a straightforward direct refusal data set. Despite its simplicity, this approach achieves safety performance on par with full-model alignment, while preserving reasoning capability close to that of the original reasoning-tuned model. This result holds for both 7B and 14B models and is validated across four benchmarks spanning mathematics, science, and code generation. It represents a rare “one stone, three birds” outcome: strong safety, strong reasoning, and computational efficiency. Moreover, we find that the performance is highly robust to the hyperparameters and configurations of LoRA.
Next, we explore the weight structure imposed by LoRA to understand the differences it introduces. We find that LoRA updates are not only low-rank by design but also exhibit smaller alignment with the weights of the original reasoning model, compared to those from full-model fine-tuning—in most layers. While the reduction in alignment is small, it may suggest that LoRA updates are less disruptive to reasoning-related weights. We further explore whether explicitly reducing such overlap—via regularization or post-hoc weight merging—can improve safety or reasoning capabilities. We observe mixed results, with one post-hoc method achieving a slightly better reasoning–safety trade-off on the AIME and GPQA datasets. More effort is still needed to develop approaches that lead to more consistent improvements across tasks, which we consider a valuable direction for future work.
@@ 4.3 Datasets and Models
Training Setup. Full-model fine-tuning is performed for 5 epochs, while LoRA fine-tuning is run for 10 epochs. We save and evaluate checkpoints at every epoch. Unless otherwise stated, LoRA is applied only to the MLP layers with rank $r=4$. In the later part of Section 4.4, we investigate the effect of varying $r$ and applying LoRA to different modules (e.g., MLP vs. attention).
@@ 4.4 LoRA is All You Need for Safety Alignment of Reasoning LLMS
For the 7B models, the best LoRA checkpoints are on par with, or even slightly outperform, the base model on three reasoning tasks (AIME, HumanEval+, and MBPP+), while showing a slight underperformance on GPQA. In terms of safety, they fall slightly short of the fully fine-tuned models, with approximately a 0.03 drop in safety score. For the 14B models, there is a persistent but small drop in reasoning accuracy compared to the base model, while the safety performance remains comparable to that of the fully fine-tuned model.
@@
(b) Safety score v.s. $r$
(c) Applying LoRA to different modules
@@
Figure 4: In (a) and (b), we show reasoning and safety performance at different LoRA ranks $r$ for the 14B model, respectively. Full-model fine-tuning is included as the rightmost point for reference. Reasoning performance decreases as $r$ increases, while safety first decreases and then increases. Overall, very low ranks (e.g., $r=1,4$) are recommended to achieve strong performance in both reasoning and safety. In (c), we compare applying LoRA to different modules. The lightness of the points indicates the number of fine-tuning epochs, with lighter colors corresponding to earlier epochs. Overall, we observe that the performance is quite similar across the two setups.
Effect of $r$. We vary the LoRA rank $r$ in the range $\{1,4,8,64\}$ for fine-tuning the 14B model, and report results at the final checkpoint in Figures 4(a) and 4(b) (with results across all epochs shown in Figure 9). We also include full-model fine-tuning as the rightmost point, since it allows for full-rank updates. From both the final checkpoint and the full training trajectories, we observe that reasoning performance generally declines as $r$ increases, although the drop is very minor between $r=1$ and $r=8$. This suggests that reasoning performance is not very sensitive to the exact choice of $r$, as long as it remains in a reasonable range. For safety scores, we observe a notable drop when $r$ increases from $4$ to $64$, whereas full-model fine-tuning yields a better safety score than $r=64$. We suspect this may be due to optimization difficulty at intermediate high ranks, while very low-rank (extremely under-parameterized) or full-rank (heavily over-parameterized) setups may benefit from easier optimization dynamics, leading to better safety outcomes. Overall, to achieve strong performance in both reasoning and safety, using a very small $r$ is recommended.
Effect of applying LoRA to different modules. Next, we examine the effect of applying LoRA to different modules. Specifically, we compare applying LoRA only to the MLP layers versus applying it to both the MLP and self-attention layers (i.e., the Q, K, V, and O projections). As shown in Figure 4(c), there is no clear difference in performance between the two setups. We note that applying LoRA only to the MLP layers is more computationally efficient.
@@ 5 Exploring the Structure of LoRA Weights
Next, we investigate the alignment between LoRA and reasoning weights and study if reducing the alignment further boosts the performance.
@@ 5.1 LoRA Updates Have Less Alignment with Initial Weights
We compare the full-rank fine-tuned model with the LoRA fine-tuned model in which both the attention and MLP modules are updated (i.e., corresponding to “QKVO & MLP” in Figure 4(c)), making the two settings more comparable since updates occur in all major modules. We compute the alignment metrics for different module types across layers, average them over layers, and report the results for the 14B models in Figure 5. We observe that LoRA achieves smaller overlap in most modules across the metrics, with a few exceptions. This suggests that, for most weights, ${\bm{W}}_{I}$ and $\Delta{\bm{W}}$ are more orthogonal—both in the column and row spaces—for LoRA than for full-model fine-tuning. In other words, under LoRA fine-tuning, the safety-oriented updates read from and write into subspaces that are more separate from those used by the original reasoning-related weights—more so than in the full-model fine-tuned version. Although the reduction in alignment values is sometimes small, it may still indicate that LoRA updates interfere less with the reasoning-related components of the model, potentially explaining the better preservation of reasoning performance. A more in-depth investigation is needed to fully understand the underlying mechanisms and to develop more precise metrics for capturing this effect—an important direction for future work.
(a) 7B models on AIME and GPQA.
(b) 7B models on HumanEval+ and MBPP+.
(c) 14B models on AIME and GPQA.
(d) 14B models on HumanEval+ and MBPP+.
Figure 6: For each method, we visualize only the points on the Pareto frontier of reasoning performance versus safety score. (a)(c) On AIME and GPQA, OrthoMerge-both appears most promising, as its points are more concentrated in the upper-right corner and are sometimes strictly better than those of vanilla LoRA—especially for the 14B models. (b)(d) However, when we compare it to vanilla LoRA on the coding benchmarks, it slightly underperforms in terms of accuracy. Overall, OrthoMerge-both can occasionally outperform vanilla LoRA in reasoning performance while achieving slightly higher safety scores.
@@ 5.2 Exploring Methods That Further Reduce Alignment
Given the observations in Section 5.1, we wondered whether further reducing the overlap between $\Delta{\bm{W}}$ and ${\bm{W}}_{I}$ could lead to even better reasoning performance without compromising safety. This question is particularly worth exploring because, while the results achieved by LoRA in Section 4.4 are strong, they are not perfect—a small performance gap remains compared to the original reasoning model, especially for the 7B model on GPQA and the 14B model on AIME, HumanEval+, and MBPP+.
To test this, we consider two approaches for enforcing stronger orthogonality: regularization during LoRA fine-tuning and enforcing orthogonality during merging LoRA weights, as we discuss next.
@@
In Figures 6(a) and 6(c), we show the reasoning performance on AIME and GPQA against safety. Since we have two evaluation dimensions—safety and reasoning performance—for each method that yields multiple checkpoints (e.g., different epochs for vanilla and regularization fine-tuning, or different hyperparameters for OrthoMerge), we only visualize the Pareto-optimal points to compare the best-performing configurations. We observe that methods addressing both the row and column spaces (“both”) tend to yield better results than those that only operate on the column space (“col”). Among them, OrthoMerge-both appears most promising, with more points concentrated in the upper-right corner—often strictly dominating the points corresponding to vanilla LoRA. Therefore, we additionally evaluate OrthoMerge-both on the coding benchmarks and compare it against vanilla in Figures 6(b) and 6(d). However, in this case, it slightly underperforms. Overall, across tasks, our conclusion is that OrthoMerge-both tends to achieve slightly better safety scores and occasionally better reasoning performance than vanilla LoRA, though the improvements are not consistent. This suggests potential for further improving the safety–reasoning trade-off by encouraging orthogonality, and further investigation is needed to identify approaches that yield consistent gains—making this a valuable direction for future research.
6 Conclusion and Discussion
In this work, we identified a simple solution to the previously observed tension between reasoning and safety in reasoning LLMs. Through extensive experiments, we found that fine-tuning a reasoning model with LoRA on a safety dataset preserves its reasoning abilities, whereas full-model fine-tuning significantly degrades them. Additionally, we observe that LoRA induces weight updates with smaller overlap with the initial weights, compared to full-model fine-tuning. This may help explain the better reasoning performance of LoRA safety fine-tuning. We also explore further reducing this overlap via regularization or during weight merging and observe slight improvements on certain tasks. We hope this can inspire future work to design methods that lead to more consistent improvements. Another promising research direction is to understand and characterize how LoRA naturally produces updates that are less disruptive to the model’s reasoning function. This is particularly intriguing because LoRA only constrains the update rank, without explicitly restricting the direction.

### v2

@@ Abstract
Reasoning LLMs have demonstrated remarkable breakthroughs in solving complex problems that were previously out of reach. To ensure LLMs do not assist with harmful requests, safety alignment fine-tuning is necessary in the post-training phase. However, safety alignment fine-tuning has recently been shown to significantly degrade reasoning abilities, a phenomenon known as the “Safety Tax”. In this work, we show that using LoRA for SFT on refusal datasets effectively aligns the model for safety without harming its reasoning capabilities. This is because restricting the safety weight updates to a low-rank space minimizes the interference with the reasoning weights. Our extensive experiments across four benchmarks covering math, science, and coding show that this approach produces highly safe LLMs—with safety levels comparable to full-model fine-tuning—without compromising their reasoning abilities. Our ablation studies further identify three key factors in LoRA: (1) rank-$1$ updates are sufficient to achieve the best reasoning and safety performance, (2) the up projection layers are the most critical modules, with LoRA applied to them alone achieving even better results, and (3) middle layers are more effective than early or late layers. Together, these findings show that strong safety and reasoning can be achieved at minimal computational cost when updates are applied in the right places. Additionally, we observe that LoRA induces weight updates with smaller overlap with the initial weights compared to full-model fine-tuning. Finally, while our attempts to further reduce this overlap yield only modest improvements on some tasks, they highlight the potential of developing methods that more reliably optimize the reasoning–safety tradeoff.
@@ 1 Introduction
Our extensive experiments reveal the surprising effectiveness of a simple recipe for safety alignment of reasoning models: Applying LoRA during SFT using a straightforward direct refusal data set. Despite its simplicity, this approach achieves safety performance on par with full-model alignment, while preserving reasoning capability close to that of the original reasoning-tuned model. This result holds for both 7B and 14B models and is validated across four benchmarks spanning mathematics, science, and code generation. It represents a rare “one stone, three birds” outcome: strong safety, strong reasoning, and computational efficiency.
Moreover, we further ablate the LoRA configuration to understand “how much LoRA is sufficient”. We make three key findings. (1) Setting the rank to $r=1$ achieves the best reasoning–safety tradeoff (in terms of the Pareto frontier when the number of training epochs is varied). This is encouraging, as it shows that strong performance on both reasoning and safety can be achieved at the lowest possible fine-tuning cost. (2) Updating only the up projection layers in the MLP yields an even better tradeoff than updating the full MLP, while updating only the gate or down projections performs worse. This highlights the central role of the up projection and motivates future research into understanding why it is so effective. (3) Middle layers are more important for a good reasoning–safety tradeoff: updating only 16 middle layers is often sufficient, whereas updating early or late layers yields worse results. Interestingly, this connects to prior findings that safety-critical features often emerge in the middle layers of LLMs (Panickssery et al., 2023; Arditi et al., 2024). Overall, our results provide meaningful insights into the key elements of LoRA configuration that affect the reasoning–safety tradeoff, and can help achieve strong performance at minimal cost.
Additionally, we explore the weight structure imposed by LoRA to understand the differences it introduces. We find that LoRA updates are not only low-rank by design but also exhibit smaller alignment with the weights of the original reasoning model compared to those from full-model fine-tuning—across most layers. While the reduction in alignment is small, it may suggest that LoRA updates are less disruptive to reasoning-related weights. We further investigate whether explicitly reducing such overlap—via regularization or post-hoc weight merging—can improve safety or reasoning capabilities. We find that one post-hoc method for reducing overlap achieves a modest gain in the reasoning–safety trade-off on some tasks. This shows promise, but more effort is needed to develop approaches that yield consistent improvements across tasks, which we consider a valuable direction for future work.
@@ 4.3 Datasets and Models
Training Setup. Full-model fine-tuning is performed for 5 epochs, while LoRA fine-tuning is run for 10 epochs. We save and evaluate checkpoints at every epoch. Unless otherwise stated, LoRA is applied only to the MLP layers with rank $r=1$. In Section 5, we investigate the effect of varying $r$ and applying LoRA to different modules (e.g., MLP vs. attention) and layers.
@@ 4.4 LoRA is All You Need for Safety Alignment of Reasoning LLMS
For the 7B models, the best LoRA checkpoints outperform the base model on all tasks. In terms of safety, they fall slightly short of the fully fine-tuned models, with an average drop of about 0.03 in safety score. For the 14B models, there is a consistent but small decrease in reasoning accuracy compared to the base model, while safety performance remains comparable to that of the fully fine-tuned model.
5 How Much LoRA is Enough?
In this section, we ablate the LoRA configuration to identify the key elements that matter most. We examine three factors: (1) the rank $r$, (2) the modules to which LoRA is applied, and (3) the layers to which LoRA is applied. Our goal is to determine the minimal setup needed for LoRA—i.e., the smallest update sufficient to achieve both strong reasoning and safety.
To compare different LoRA configurations, we visualize performance in the reasoning–safety plane. For each configuration, we train for 10 epochs, which yields multiple checkpoints with varying reasoning and safety scores. We then evaluate configurations by examining the Pareto frontier of these checkpoints, which allows us to compare tradeoffs effectively.
5.1 Rank: $r=1$ is Sufficient
We explore the effect of the rank $r$ in LoRA. Specifically, we run experiments with the 14B model by applying LoRA to the MLP layers while varying $r$. Results at the final checkpoint are shown in Figures 4(a) and 4(b). We also include full-model fine-tuning as the rightmost point, since it allows for full-rank updates. Reasoning performance generally declines as $r$ increases, although the drop is minor between $r=1$ and $r=8$. For safety scores, we observe a notable decrease when $r$ increases from 4 to 64, whereas full-model fine-tuning yields a better safety score than $r=64$. We suspect this may be due to optimization difficulties at intermediate high ranks, while very low-rank (under-parameterized) or full-rank (over-parameterized) setups may benefit from easier optimization dynamics, leading to better safety outcomes. Most importantly, we find that $r=1$ is sufficient to achieve the best performance in both reasoning and safety. Figure 4(c) further confirms that $r=1$ also achieves the best (on AIME) or nearly best (on GPQA) tradeoff when varying the number of training epochs. This indicates that we can achieve strong performance on both reasoning and safety at the lowest fine-tuning cost.
Additionally, the fact that $r=1$ is sufficient reflects the inherently low-rank nature of the safety alignment task itself. This connects to prior work showing that safety can be mediated by a single direction within the model’s internal representations, often referred to as a steering vector (Panickssery et al., 2023) or refusal features (Arditi et al., 2024; Yu et al., 2024). This perspective may explain why a rank-1 update is enough to achieve safety.
@@ 5.1 Rank: $r=1$ is Sufficient
(b) Safety v.s. $r$
(c) Reasoning-safety tradeoff
@@
Figure 4: In (a) and (b), we show reasoning and safety performance at different LoRA ranks $r$ for the 14B model, respectively. Full-model fine-tuning is included as the rightmost point for reference. Reasoning performance decreases as $r$ increases, while safety first decreases and then increases. Overall, very low ranks are recommended, and $r=1$ already achieves the best performance in both metrics. In (c), we visualize the Pareto frontiers of the reasoning–safety tradeoff when the training epoch is varied and observe that $r=1$ is sufficient to achieve an excellent tradeoff, outperforming other ranks, especially on AIME.
5.2 Modules: Up-Projection Matters Most in MLPs
Figure 5: Applying LoRA to MLP modules alone is sufficient. Faded points indicate non–Pareto-frontier points.
The LoRA adapter is usually applied to attention layers and/or MLP modules, e.g., by setting target_modules when using the PEFT package. Here, we explore the effect of applying LoRA to different modules. We first compare applying it to both attention and MLP layers (QKVO & MLP in the figure) versus only applying it to MLP layers. Figure 5 shows the comparison for the 14B models with $r=1$. We observe that applying LoRA only to the MLP layers yields a similar Pareto frontier compared to applying it to both modules.
Next, we perform an ablation over the modules within the MLP. Specifically, in the Qwen architecture, the MLP layers use the popular SwiGLU Chowdhery et al. (2022) structure that contains a gate projection, an up projection, and a down projection. We apply LoRA to only one of these projection layers at a time and evaluate the results. We use $r=1$. Figure 6 shows that across different tasks, applying LoRA only to the up projection achieves strong results, with the Pareto frontier often on par with using the full MLP and even outperforming it on the two coding benchmarks. In contrast, applying LoRA only to the down projection yields noticeably worse performance. These findings suggest that different projections within the MLP contribute differently to the reasoning–safety tradeoff, and that the up projection is particularly important and sufficient by itself in our setup.
Figure 6: We compare applying LoRA to different projections within the MLP layers. The results show that applying it only to the up projection achieves the best tradeoff, and even outperforms applying it to the full MLP on the coding benchmarks.
Discussion. We mainly focus our ablations on the MLP, as it is primarily responsible for feature transformations and thus well-suited for safety alignment. However, we believe it is also worthwhile to study ablations over the attention layers, which we leave for future work. It is also valuable to further investigate why the up projection alone is the best choice. Additionally, extending these experiments to other LLM architectures would be an important direction for future work.
5.3 Layers: Middle Layers Matter Most
Figure 7: Applying LoRA to the middle layers (17–32) achieves a better tradeoff compared to using either the early or late layers, and performs on par with or only slightly behind using all layers across tasks. In the plot, the faded points indicate non–Pareto-frontier points.
We ablate over layers in the model. The 14B Qwen2.5 architecture has 48 layers in total, and we apply LoRA to only 16 of them. We consider three configurations: (1) layers 5–20, denoted as “Early Layers”, (2) layers 17–32, denoted as “Middle Layers”, and (3) layers 25–40, denoted as “Late Layers”. In all cases, we apply LoRA only to the up projection layers with $r=1$. The results are shown in Figure 7. Across tasks, we observe that applying LoRA to the middle layers yields the best tradeoff, achieving performance on par with using all layers on AIME and GPQA, and only slightly behind using all layers on HumanEval+ and MBPP+. In contrast, applying LoRA to either the early or late layers results in a noticeably worse tradeoff. This shows that the middle layers are most important for balancing reasoning and safety. Interestingly, this again connects to prior findings on steering vectors (Panickssery et al., 2023) and refusal features (Arditi et al., 2024; Yu et al., 2024), which suggest that directions in intermediate representations responsible for safety behavior are most prominent in the middle layers.
@@ 6.1 LoRA Updates Have Less Alignment with Initial Weights
We compare the full-rank fine-tuned model with the LoRA fine-tuned model in which both the attention and MLP modules are updated with $r=4$, making the two settings more comparable since updates occur in all major modules. We compute the alignment metrics for different module types across layers, average them over layers, and report the results for the 14B models in Figure 8. We observe that LoRA achieves smaller overlap in most modules across the metrics, with a few exceptions. This suggests that, for most weights, ${\bm{W}}_{I}$ and $\Delta{\bm{W}}$ are more orthogonal—both in the column and row spaces—for LoRA than for full-model fine-tuning. In other words, under LoRA fine-tuning, the safety-oriented updates read from and write into subspaces that are more separate from those used by the original reasoning-related weights—more so than in the full-model fine-tuned version. Although the reduction in alignment values is sometimes small, it may still indicate that LoRA updates interfere less with the reasoning-related components of the model, potentially explaining the better preservation of reasoning performance. A more in-depth investigation is needed to fully understand the underlying mechanisms and to develop more precise metrics for capturing this effect—an important direction for future work.
(a) 14B models on AIME and GPQA.
(b) 14B models on HumanEval+ and MBPP+.
Figure 9: For each method that produces multiple checkpoints (e.g., across epochs or hyperparameter settings), we visualize the Pareto-frontier points. Methods enforcing orthogonality in both row and column spaces perform better than column-only variants. The post-hoc method OrthoMerge-both is the most promising, with points concentrated in the upper-right corner and its best point strictly dominating vanilla LoRA on AIME and GPQA. On MBPP+, it achieves a slightly better Pareto frontier, while on HumanEval+ it slightly underperforms.
@@ 6.2 Exploring Methods That Further Reduce Alignment
Given the observations in Section 6.1, we ask whether further reducing the overlap between $\Delta{\bm{W}}$ and ${\bm{W}}_{I}$ could lead to even better reasoning performance without compromising safety. This question is particularly relevant because, while the results achieved by LoRA in Section 4.4 are strong, they are not perfect—a small performance gap remains compared to the original reasoning model, especially for the 14B model on AIME, HumanEval+, and MBPP+.
We experiment with two approaches: (1) Regularization during LoRA training: adding a penalty to discourage overlap between ${\bm{W}}_{I}$ and $\Delta{\bm{W}}$, targeting either the column space (reg-col) or both the column and row spaces (reg-both). For efficiency, we approximate ${\bm{W}}_{I}$ via a truncated SVD. We tried different values of $\beta$ but observed negligible differences, so we fix $\beta=1$. (2) Post-hoc Orthogonalization: post-processing $\Delta{\bm{W}}$ before applying Equation 1, projecting it onto the orthogonal complement of the top-$k$ singular vectors of ${\bm{W}}_{I}$. We test both column-only (OrthoMerge-col) and column+row (OrthoMerge-both) variants, with scaling $\lambda$ to offset safety loss. Further details are deferred to Appendix B.
@@
Figure 9(a) shows the results on AIME and GPQA for the 14B models with $r=4$. For each method that yields multiple checkpoints (e.g., different epochs for vanilla and regularization fine-tuning, or different hyperparameters for OrthoMerge), we visualize the Pareto-frontier points. We observe that methods addressing both the row and column spaces (“both”) tend to yield better results than those that only operate on the column space (“col”). Among them, OrthoMerge-both appears most promising, with more points concentrated in the upper-right corner and its best point strictly dominating vanilla LoRA. Therefore, we additionally evaluate OrthoMerge-both on the coding benchmarks in Figure 9(b), where it slightly underperforms on HumanEval+ but slightly outperforms vanilla on MBPP+ in terms of the tradeoff.
Overall, we observe modest yet inconsistent gains from post-hoc orthogonalization. This highlights the potential of controlling the subspace geometry of LoRA updates and points to the need for more nuanced methods that yield consistent improvements.
7 Conclusion
In this work, we identified a simple yet effective solution to the previously observed tension between reasoning and safety in LLMs. Through extensive experiments, we showed that applying LoRA during safety fine-tuning preserves reasoning capabilities while achieving strong safety alignment, in contrast to full-model fine-tuning, which significantly degrades reasoning. We further ablate the LoRA configuration to identify the key elements and make three important findings: (1) rank-$1$ updates are sufficient to achieve the best reasoning–safety tradeoff, (2) the up projection modules are most important for the tradeoff, yielding even better results than updating all or other modules, and (3) middle layers matter most, more than early or late layers. These results highlight that strong safety and reasoning can be achieved at minimal computational cost when updates are applied in the right places. Finally, we analyze the weight structure of LoRA and find that its updates are more orthogonal to the initial weights, which may help explain their reduced interference with reasoning. While our attempts to further enforce orthogonality yield only modest and inconsistent improvements, they highlight the potential of better controlling the geometry of safety updates as an avenue for future research. Overall, our results provide both a practical recipe for safety alignment of reasoning models and new insights into why LoRA is uniquely well-suited to this problem.
@@ A.5 Training Details
Appendix B Detailed Description of the Methods Used in Section 6.2
We consider two approaches for enforcing stronger orthogonality: regularization during LoRA fine-tuning and enforcing orthogonality during merging LoRA weights, as we discuss next.
Appendix C Additional Figures
Figure 11 shows the results on the base versions of HumanEval and MBPP.